OpenStudy (zubhanwc3):

Intergration help, 2 questions

OpenStudy (zubhanwc3):

anyone that cant view the formula use this http://openstudy.com/study#/updates/52de57e0e4b003c643a10b24 answer the following problems by integrating exponential functions in bases other than e: 1)\[\int\limits_{}^{}4^{^{cosx}}\sin(x)dx\]2)\[\int\limits_{}^{}10^{x}dx\]